# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ 4. The CM case and magnetic modular forms
We will now specialize to the CM case. Let $D<0$ (with $D\neq-3,-4$) be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point with discriminant $D$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by $\mathcal{O}_{D}$ so that $j(C)=j(\alpha_{D})$ and let $\psi_{C}$ be the corresponding Hecke character of $K:=\mathbb{Q}(\sqrt{D})$. For the convenience of the later discussion, we will choose $C/\mathbb{Q}$ in a way that all of its bad primes divide $D$. By Deuring’s criterion, the following are equivalent for a prime $p$ with $p\nmid D$:
@@
    $p$ is a supersingular prime of $C$ (resp. $p$ is an ordinary prime of $C$);
@@
    $p$ is inert in $\mathcal{O}_{D}$ (resp. $p$ splits in $\mathcal{O}_{D}$);
@@
    $(\frac{D}{p})=-1$ (resp. $(\frac{D}{p})=1$).
@@
Now, on the elliptic curve side, we have the following splitting (as Galois representations of $K$)
(7) $\mathrm{Res}_{K/\mathbb{Q}}(\mathrm{Sym}^{k-2}C)\;\cong\;\mathrm{Res}_{K/\mathbb{Q}}\big(\mathrm{Sym}^{k-2}(\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C})\big)\;\cong\;\bigoplus_{\begin{subarray}{c}a+b=k-2\\ 0\leq a,b\leq k-2\end{subarray}}\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}.$
Let $p$ be a good prime of $C$. Then,
@@
    if $p=\mathfrak{p}$ is inert in $\mathcal{O}_{D}$, then $\mathrm{Frob}_{\mathfrak{p}}(=\mathrm{Frob}_{p}^{2})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{k-2}$;
@@
    if $p$ splits in $\mathcal{O}_{D}$, then for any embedding $\pi:K\hookrightarrow\mathbb{Q}_{p}$, $\mathrm{Frob}_{\pi}(=\mathrm{Frob}_{p})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $u_{p}(C)^{a-b}p^{b}$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit under the embedding $\pi$. [^4] Technically, $\mathrm{Frob}_{\pi}$ acts on one of $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}{\hskip 1.0pt}\overline{\psi_{C}^{a}}$ by $u_{p}(C)^{a-b}p^{b}$ depending on the choice of $\psi_{C}$. We will thus replace $\psi_{C}$ by $\overline{\psi_{C}}$ if necessary to ensure that $\mathrm{Frob}_{\pi}$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $u_{p}(C)^{a-b}p^{b}$.
This suggests that on the modular form side, the space
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(\alpha_{D})},\;\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\bigg\}$
should also split into $1$-dimensional pieces $\{V_{a,b}\}_{a+b=k-2}$ equipped with corresponding $U_{p}$ actions. In other words, there should exist $\{G_{a,b}\}_{a+b=k-2}$ such that
  1. (1)
    $\mathrm{Span}\{G_{a,b}\mid a+b=k-2\}=\mathrm{Span}\big\{\frac{E_{k}}{j-j(\alpha_{D})},\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\big\}$;
  2. (2)
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=-1$, $U_{p}^{2}G_{a,b}\equiv p^{k-2}G_{a,b}$ modulo some power of $p$;
  3. (3)
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=1$, $U_{p}G_{a,b}\equiv u_{p}(C)^{a-b}p^{b}G_{a,b}$ modulo some power of $p$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
We will now describe the (numerical) construction of these $G_{a,b}$.
Recall that for a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$, we may view $f$ as a function in $\tau$ and $\overline{\tau}$ and define the nonholomorphic derivative as
@@
For $k\in\{4,6,8,10,14\}$, write  [^5] This two-variable modular form $G_{k}(z,\tau)$ occurred in many places throughout the literature. We refer to [AKN97, p. 98] for a detailed discussion and the references therein. It (and its half-integral weight analogue) also occurred in [Zag02, DJ08a] and plays a key role in Zagier’s duality.
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)},$
and for $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;(\partial_{\tau}^{r-1}G_{k})(z,\alpha_{D}).$
While $G_{k,D}^{(r)}$ a priori depends on the choice of $\alpha_{D}$ and may even be identically zero, we will describe a way to produce a reasonably standard candidate, which we also denote by $G_{k,D}^{(r)}$.
  1. (1)
    If $G_{k,D}^{(r)}$ is identically zero  [^6] In practice, it seems that if $D\neq-3,-4$, then $G_{k,D}^{(r)}$ is never identically zero and in particular, $a_{1}(G_{k,D}^{(r)})\neq 0$., then there should exist some $s\in\mathbb{Z}^{+}$ such that $\frac{(\partial_{\tau}^{r-1}G_{k})(z,\tau)}{(j(\tau)-j(\alpha_{D}))^{s}}$ is not identically zero at $\tau=\alpha_{D}$. We may thus replace $G_{k,D}^{(r)}$ by $\frac{(\partial_{\tau}^{r-1}G_{k})(z,\tau)}{(j(\tau)-j(\alpha_{D}))^{s}}\big|_{\tau=\alpha_{D}}$.
  2. (2)
    By the CM theory, there should exist some nonzero $A=A_{k,D}^{(r)}\in\mathbb{C}$ such that $AG_{k,D}^{(r)}$ has integer coefficients. We may choose $A$ appropriately and replace $G_{k,D}^{(r)}$ by $AG_{k,D}^{(r)}$ so that the coefficients of $G_{k,D}^{(r)}$ have no common divisors and that the first nonzero coefficient of $G_{k,D}^{(r)}$ is positive.
The downside of making a standard choice of $G_{k,D}^{(r)}$ is that every phenomena stated for $G_{k,D}^{(r)}$ in this section may only hold up to a nonzero integer multiple.
Now, each $G_{k,D}^{(r)}$ has a pole at $\alpha_{D}$ of order exactly $r$ and is of the form
$A_{1}\cdot\frac{E_{k}}{j-j(\alpha_{D})}+A_{2}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{2}}+\cdots+A_{r}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{r}}$
for some $A_{i}\in\mathbb{Z}$. Through the following conjectures, one observes that each $G_{k,D}^{(r)}$ takes the role of $G_{k-1-r,r-1}$. Here the supersingular case is simply a restatement of 3.3 and the ordinary case should be viewed as a natural generalization of 2.4 in the CM case.
Conjecture 4.1.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$ and let $C/\mathbb{Q}$ be an elliptic curve with $j(C)=j(\alpha_{D})$. Let $p$ be a prime with $(\frac{D}{p})=-1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;p^{k-2}a_{np^{l-2}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l-r}}$.
@@
As with the case of 3.3, it is also possible to deduce 3.1 from 4.2 in the ordinary and CM case through involved computations which we omit here.
As mentioned in the introduction, the magnetic property shows up in all the CM cases. Specifically, we have the following conjecture.
@@
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Then, $G_{k,D}^{(r)}$ is $(r^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(G_{k,D}^{(r)}).$

### v2

@@ 4. The CM case, supercongruences, and magnetic modular forms
We will now specialize to the CM case. Let D<4D<-4 be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, $\alpha_{D}\in\mathcal{H}$ be a fixed CM point with discriminant $D$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by $\mathcal{O}_{D}$ so that $j(C)=j(\alpha_{D})$ and let $\psi_{C}$ be the corresponding Hecke character of $K:=\mathbb{Q}(\sqrt{D})$. For the convenience of the later discussion, we will choose $C/\mathbb{Q}$ in a way that all of its bad primes divide $D$. By Deuring’s criterion, the following are equivalent for a prime $p$ with $p\nmid D$:
@@
    $p$ is an ordinary prime of $C$ (resp. $p$ is a supersingular prime of $C$);
@@
    $p$ splits in $\mathcal{O}_{D}$ (resp. $p$ is inert in $\mathcal{O}_{D}$);
@@
    $(\frac{D}{p})=1$ (resp. $(\frac{D}{p})=-1$).
@@
4.1. A “canonical” basis
As mentioned briefly in the Section 1.3, we will construct a basis for the space
$\mathrm{MMF}_{k,D}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(\alpha_{D})},\;\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\bigg\}$ (7)
that is diagonal (resp. anti-diagonal) with respect to the action of (powers of) $U_{p}$ via explicit supercongruences for all ordinary (resp. supersingular) primes $p$, where the (anti-)eigenvalues will be described explicitly using the $p$-adic analogue of the Chowla–Selberg periods.
To motivate the construction, we first focus on the ordinary case. Note that on the elliptic curve side, we have the following splitting (as Galois representations of $K$)
$\mathrm{Res}_{K/\mathbb{Q}}(\mathrm{Sym}^{k-2}C)\;\cong\;\mathrm{Res}_{K/\mathbb{Q}}\big(\mathrm{Sym}^{k-2}(\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C})\big)\;\cong\;\bigoplus_{\begin{subarray}{c}a+b=k-2\\ 0\leq a,b\leq k-2\end{subarray}}\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}.$ (8)
Let $p$ be a prime that splits in $\mathcal{O}_{D}$. Then, for any embedding $\pi:K\hookrightarrow\mathbb{Q}_{p}$, $\mathrm{Frob}_{\pi}(=\mathrm{Frob}_{p})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{b}u_{p}(C)^{a-b}$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit under the embedding $\pi$. [^10] Technically, $\mathrm{Frob}_{\pi}$ acts on one of $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}{\hskip 1.0pt}\overline{\psi_{C}^{a}}$ by $p^{b}u_{p}(C)^{a-b}$ depending on the choice of $\psi_{C}$. We will thus replace $\psi_{C}$ by $\overline{\psi_{C}}$ if necessary to ensure that $\mathrm{Frob}_{\pi}$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{b}u_{p}(C)^{a-b}$. This suggests that on the modular form side, the space $\mathrm{MMF}_{k,D}$ should also split into $1$-dimensional pieces $\{V_{a,b}\}_{a+b=k-2}$ equipped with corresponding $U_{p}$-actions. In other words, there should exist $\{G_{a,b}\}_{a+b=k-2}$ such that
@@ 4.1. A “canonical” basis
    $\mathrm{Span}\{G_{a,b}\mid a+b=k-2\}=\mathrm{MMF}_{k,D}$;
@@
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=1$, $U_{p}G_{a,b}\equiv p^{b}u_{p}(C)^{a-b}G_{a,b}$ modulo some power of $p$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
In order to construct these $G_{a,b}$, the idea is to view the original $\frac{E_{k}}{j-j(C)}$ as a two variable function $\frac{E_{k}(z)}{j(z)-j(\tau)}$, multiply it by some auxiliary function $f^{\mathrm{aux}}(\tau)$, differentiate the product (several times) with respect to the auxiliary variable $\tau$, and evaluate $\tau$ at the CM point $\alpha_{D}$.
In this case, the choice of the auxiliary function was already revealed in Section 1.2. That is, for $k\in\{4,6,8,10,14\}$, we choose the auxiliary function as $f_{k}^{\mathrm{aux}}(\tau)=\frac{E_{14-k}}{\Delta}(\tau)$ and consider
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)}{j(z)-j(\tau)}\cdot f_{k}^{\mathrm{aux}}(\tau)\;=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}.$ (9)
We also recall from Section 1.2 that the non-holomorphic derivative of a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$ is defined as
@@
Now, if one naively considers the derivative $(\partial_{\tau}^{r}G_{k})(z,\tau)$ and evaluates $\tau$ at $\alpha_{D}$, then the resulting modular form will depend on the choice of $\alpha_{D}$, since $(\partial_{\tau}^{r}G_{k})(z,\tau)$ is of weight 2k+2r2-k+2r with respect to $\tau$. In order to remove the dependency, we adopt the following normalization.
For $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;\bigg(\frac{\Delta}{E_{14-k}}(\tau)\cdot\bigg(-\frac{E_{4}}{E_{6}}(\tau)\bigg)^{r-1}\cdot(\partial_{\tau}^{r-1}G_{k})(z,\tau)\bigg)\bigg|_{\tau=\alpha_{D}}.$ (10)
In this way, $G_{k,D}^{(r)}$ is independent of the choice of $\alpha_{D}$, has rational coefficients  [^11] The rationality is because $D$ is assumed to have class number $1$ (see also Example 4.1 and Example 4.3)., and is of the form
$G_{k,D}^{(r)}\;=\;\sum_{i=1}^{r-1}A_{i}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{i}}+(r-1)!\cdot j(\alpha_{D})^{r-1}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{r}}.$ (11)
Through the following conjecture, one observes that each $G_{k,D}^{(r)}$ takes the role of $G_{k-1-r,r-1}$. The conjecture should also be viewed as a natural generalization of 2.2 in the CM case.
@@
It is possible to rewrite this purely in terms of negative discriminants as in 2.8. Since $(\frac{D}{p})=1$, we have that $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;\pi^{k-1-r}{\hskip 1.0pt}\overline{\pi}^{r-1}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{\overline{\pi}^{(k-1)l}}$.
Remark 4.3.
As with the case of 3.4, it is also possible to deduce the ASD congruence (3.1) in the ordinary and CM case from 4.1 through involved computations which we omit here.
Remark 4.4.
One can also “explain” 3.4 in this CM case through the same heuristic. If $p=\mathfrak{p}$ is inert in $\mathcal{O}_{D}$, then $\mathrm{Frob}_{\mathfrak{p}}(=\mathrm{Frob}_{p}^{2})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{k-2}$. Hence, for the potential basis $\{G_{a,b}\}_{a+b=k-2}$, one would expect that $U_{p}^{2}G_{a,b}\equiv p^{k-2}G_{a,b}$ modulo some power of $p$ for supersingular primes $p$, as observed in 3.4.
Now we turn to the supersingular primes. In this case, we can already inherit certain property from 3.4. However, through the help of the $p$-adic Gamma function, one can moreover observe that for a supersingular prime $p$, the family $\{G_{k,D}^{(r)}\}$ behaves like an anti-diagonal basis with respect to the $U_{p}$ operator, i.e., $U_{p}G_{k,D}^{(r)}$ is a multiple of $G_{k,D}^{(k-r)}$ modulo some power of $p$. Indeed, such phenomenon was already hinted at in [ALS25, Corollary 1.2 and Example 3].
For a fundamental discriminant $D<0$ and a prime $p$ with $p\nmid D$, define
$\Omega_{p,D}\;:=\;-\bigg(\prod_{a=1}^{|D|-1}\Gamma_{p}\bigg(\frac{a}{|D|}\bigg)^{(\frac{D}{a})}\bigg)^{\frac{w_{D}}{2}}\;\in\;\mathbb{Z}_{p}^{\times},$ (12)
where $\Gamma_{p}$ denotes the $p$-adic Gamma function, $(\frac{D}{a})$ denotes the Kronecker symbol, and $w_{D}=|\mathcal{O}_{D}^{\times}|$. One should view this definition as a $p$-adic analogue of (the square of) the Chowla–Selberg period [CS67] (see also Remark 4.10).
@@
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let $D<0$ be a discriminant with class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Suppose that $D\notin\{-3,-4,-12\}$. Let $p$ be a prime with $(\frac{D}{p})=-1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$(r-1)!\cdot a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;(k-1-r)!{\hskip 1.0pt}c_{D}^{\frac{k}{2}-r}\cdot(-p)^{r-1}\Omega_{p,D_{0}}^{-(\frac{k}{2}-r)}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$,
where we write $j_{D}=j(\alpha_{D})$ and
$c_{D}\;=\;\begin{cases}(-1)^{D}\cdot|j_{D}|^{2/3}\cdot(j_{D}-1728)^{-1}\cdot D\;(\in\mathbb{Q})&\text{if }D\neq-16,-27,\\ (-1)^{D}\cdot|j_{D}|^{2/3}\cdot\big(2(j_{D}-1728)\big)^{-1}\cdot D\;=\;-\frac{484}{3969}&\text{if }D=-16,\\ (-1)^{D}\cdot|\frac{j_{D}}{3}|^{2/3}\cdot(j_{D}-1728)^{-1}\cdot D\;=\;-\frac{3600}{64009}&\text{if }D=-27.\end{cases}$
@@
The value of $c_{D}$ is in fact always a rational square in all the cases we consider here.
$D$ 7-7 8-8 11-11 16-16 19-19 27-27 28-28 43-43 67-67 163-163
$c_{D}$ $(5/9)^{2}$ $(5/7)^{2}$ $(4/7)^{2}$ $(22/63)^{2}$ $(4/9)^{2}$ $(60/253)^{2}$ $(170/513)^{2}$ $(40/189)^{2}$ $(220/1953)^{2}$ $(26680/1672209)^{2}$
Table 1. The value of $c_{D}$.
It is also possible to describe the $p$-adic behaviors of $\{G_{k,D}^{(r)}\}$ for ordinary primes $p$ using the $p$-adic Gamma function, though we would need to exclude the case when D=12D=-12 as before.
Conjecture 4.7.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let $D<0$ be a discriminant with class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Suppose that $D\notin\{-3,-4,-12\}$. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;p^{r-1}\Omega_{p,D_{0}}^{\frac{k}{2}-r}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$.
Remark 4.8.
If one renormalizes $G_{k,D}^{(r)}$ by setting
$\widehat{G}_{k,D}^{(r)}\;:=\;(k-1-r)!c_{D}^{-\frac{r-1}{2}}\cdot G_{k,D}^{(r)}$
and renumbers them by setting $H_{k,D}^{(i)}:=\widehat{G}_{k,D}^{(\frac{k}{2}+i)}$ for $|i|\leq\frac{k}{2}-1$, then one can write 4.5 and 4.7 in the following uniform way: for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(H_{k,D}^{(\epsilon_{p}i)})\;\equiv\;\big(\epsilon_{p}p)^{\frac{k}{2}+i-1}\Omega_{p,D_{0}}^{-\epsilon_{p}i}{\hskip 1.0pt}a_{np^{l-1}}(H_{k,D}^{(i)})\pmod{p^{(k-1)l}}$,
where $\epsilon_{p}=(\frac{D}{p})$. While the issue here is that $c_{D}$ is not a square, a naive workaround is to simply view the congruence as describing that the difference of both sides has $p$-adic valuation $\geq(k-1)l$.
Remark 4.9.
Let $D<0$ be a fundamental discriminant with class number $1$. By a result of Gross–Koblitz [GK79, Theorem 4.13], one can give an explicit characterization of $\Omega_{p,D}$ when $(\frac{D}{p})=1$. Suppose that $p$ splits as $p=\pi\overline{\pi}$ in $\mathcal{O}_{D}$ and let $\iota_{\pi}:\mathcal{O}_{D}\rightarrow\mathbb{Z}_{p}$ be the $p$-adic embedding induced by $\pi$. Then,
$\Omega_{p,D}\;=\;\iota_{\pi}(\overline{\pi}^{2}).$
Equivalently, if one chooses an elliptic curve $C/\mathbb{Q}$ with CM by $\mathcal{O}_{D}$ in a way that all of its bad primes divide $D$, then
$\Omega_{p,D}\;=\;u_{p}(C)^{2},$
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit. In this way, 4.7 is equivalent to 4.1 when $D$ is fundamental.
When $D$ is non-fundamental, i.e., D=16,27,28D=-16,-27,-28 in our case, one observes that the elliptic curve $C_{D}/\mathbb{Q}$ with CM by $\mathcal{O}_{D}$ such that all of its bad primes divide $D$ is isogenous to the one $C_{D_{0}}/\mathbb{Q}$ with CM by $\mathcal{O}_{D_{0}}$ such that all of its bad primes divide $D_{0}$, where $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. In this way, it follows that
$\Omega_{p,D_{0}}\;=\;u_{p}(C_{D_{0}})^{2}\;=\;u_{p}(C_{D})^{2}.$
Hence, 4.7 is also equivalent to 4.1 in the non-fundamental case.
Remark 4.10.
One can also relate $\Omega_{p,D}$ with the special value of a $p$-adic $L$-function by a result of Ferrero–Greenberg [FG79, Theorem 1]. Let $D<0$ be a fundamental discriminant and fix a prime $p$ with $p\nmid D$. Let $\chi=(\frac{D}{\cdot})$ be the (primitive) Dirichlet character associated to $D$ and let $\varepsilon_{p}$ denote the $p$-adic cyclotomic character. Then,
$L_{p}^{\prime}(0,\chi\varepsilon_{p})\;=\;\sum_{a=1}^{|D|-1}\chi(a)\log_{p}\Gamma_{p}\bigg(\frac{a}{|D|}\bigg)-L_{p}(0,\chi\varepsilon_{p})\cdot\log_{p}|D|,$ (13)
and hence
$L_{p}^{\prime}(0,\chi\varepsilon_{p})\;=\;\frac{2}{w_{D}}\log_{p}\Omega_{p,D}-L_{p}(0,\chi\varepsilon_{p})\cdot\log_{p}|D|,$
where $L_{p}$ denotes the $p$-adic $L$-function and $\log_{p}$ denotes the $p$-adic logarithm. Here $L_{p}(0,\chi\varepsilon_{p})$ can be expressed explicitly as the following
$L_{p}(0,\chi\varepsilon_{p})\;=\;-\big(1-\chi(p)\big)\cdot\frac{1}{|D|}\sum_{a=1}^{|D|-1}\chi(a)a.$
The formula (13) should be compared with the version of the complex $L$-functions. In this case, Lerch’s formula [Ler97, p. 303, (25)] says that
$L^{\prime}(0,\chi)$ $\;=\;\sum_{a=1}^{|D|-1}\chi(a)\log\Gamma\bigg(\frac{a}{|D|}\bigg)-L(0,\chi)\cdot\log|D|.$
4.2. The magnetic property and other observations
In fact, one can already observe interesting numerical phenomena for the family $\{G_{k,D}^{(r)}\}$ by viewing $\mathrm{Sym}^{k-2}C$ simply as a Galois representation of $\mathbb{Q}$. Indeed, we have the following splitting (as Galois representations of $\mathbb{Q}$) [^12] A possibly easier way of observing this splitting is through the splitting of the $L$-function of $\mathrm{Sym}^{k-2}C$.
$\mathrm{Sym}^{k-2}C\;\cong\;\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}\oplus{\hskip 1.0pt}\bigoplus_{r=1}^{\frac{k-2}{2}}\big(\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}\big),$ (14)
where $(\frac{D}{\cdot})$ denotes (the Dirichlet character given by) the Kronecker symbol and $\varepsilon_{\ell}$ denotes the $\ell$-adic cyclotomic character. Here the term $\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}$ should correspond to $G_{k,D}^{(k/2)}$ and the term $\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}$ should correspond to $\mathrm{Span}\{G_{k,D}^{(r)},G_{k,D}^{(k-r)}\}$. We also observe that the appearance of powers of $\varepsilon_{\ell}$ seems to account for the magnetic property of $G_{k,D}^{(r)}$. Specifically, we have the following.
Conjecture 4.11.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$. Let $A=A_{k,D}^{(r)}$ be a nonzero integer such that $\widetilde{G}_{k,D}^{(r)}=AG_{k,D}^{(r)}$ has integer coefficients with no common divisors. Then, $\widetilde{G}_{k,D}^{(r)}$ is $(r^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{G}_{k,D}^{(r)}),$
where $r^{\prime}=\min(r,k-r)$.
Remark 4.12.